Help solve this urgently please:
2a^2-2ab+b^2 (when a= 3 and b= -1)
THANK YOU!!!!!!

Answer by nerdybill(7384)

2a^2-2ab+b^2 (when a= 3 and b= -1)
.
Simply plug in the given values:
2(3)^2-2(3)(-1)+(-1)^2
= 2(9)-2(3)(-1)+1 (take care of exponents first)
= 18-2(-3)+1 (multiplication next)
= 18+6+1
= 25